OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 4 x - 2y = -5 2 -2 1 -1

jimthompson5910 (jim_thompson5910):

solve the 2nd equation for x x - 2y = -5 x = -5 + 2y

jimthompson5910 (jim_thompson5910):

then plug this into the first equation 2x + 3y = 4 2(-5 + 2y) + 3y = 4 -10 + 4y + 3y = 4 -10 + 7y = 4 I'll let you finish

OpenStudy (anonymous):

plux in first equation 2(-5+2y) +3y=4

jimthompson5910 (jim_thompson5910):

solve for y, then use the solution for y to get x

OpenStudy (anonymous):

7y=14 so y=2 just replace it you find you answer

OpenStudy (anonymous):

Its -1! Lol thanks guys

jimthompson5910 (jim_thompson5910):

yep, you got it x = -5 + 2y x = -5 + 2(2) x = -5 + 4 x = -1

OpenStudy (anonymous):

:D